# Data $\times$ LLM: From Principles to Practices

### v2

@@ 2.4.2 Data Compression
(1) Metric-Based Compression. Some studies [191, 192], based on the hypothesis that a vast amount of knowledge is stored in the model parameters, have proposed methods to compress prompts while minimizing information loss. LLMLingua [191] uses a self-information criterion to remove redundant tokens from the original prompt. By quantifying the negative logarithmic probability (information entropy) of each token, LLMLingua identifies and removes tokens that can be predicted from the model’s inherent knowledge, thereby shortening the prompt while retaining essential context.
@@
(2) Finetuned-Model-Based Compression. Some methods [305, 295, 103] use fine-tuned models to compress prompts. Unlike direct pruning strategies, LLMLingua-2 [305] defines prompt compression as a problem of classifying tokens and trains a dedicated model for compression. It uses a Transformer encoder to capture bidirectional contextual information, ensuring that the compressed prompt is faithful to the original. [295] proposes a technique called ’gisting’, where a language model is trained to condense the prompt into a compact ’gist token’. These tokens encapsulate the core semantic content of the prompt and can be cached for later use. This method achieves a compression rate of up to 26 times.  [103] suggests a method to transform pre-trained language models into AutoCompressors. The AutoCompressor compresses long contexts into summary vectors, and training is performed on the model parameters using these summary vectors.
@@ 2.4.3 Data Packing
Short Sequence Insertion. Some methods [117, 261] involve inserting short sequences into long sequences to minimize padding. The Best-fit Packing method [117] first splits long documents according to the model’s context length, then sorts all document blocks in descending order of length. For each document block, it selects the training sequence set with the smallest remaining capacity that can accommodate it. [261] prioritizes long documents and uses a greedy algorithm to fill remaining space with short document segments (sequences), reducing padding and minimizing document concatenation to lower contextual noise.

### v3

@@ 2.5.2 Data Compression
(1) Metric-Based Compression. Some studies [189, 190], based on the hypothesis that a vast amount of knowledge is stored in the model parameters, have proposed methods to compress prompts while minimizing information loss. LLMLingua [189] uses a perplexity criterion to remove redundant tokens from the original prompt. By quantifying the negative logarithmic probability (perplexity) of each token through a small model, LLMLingua identifies and removes tokens that can be predicted from the model’s inherent knowledge, thereby shortening the prompt while retaining essential context.
@@
(2) Finetuned-Model-Based Compression. Unlike the aforementioned methods that use a small model’s perplexity for compression, some methods [303, 293, 102] directly perform the compression task end-to-end by fine-tuning a model. LLMLingua-2 [303] defines prompt compression as a problem of classifying tokens and trains a dedicated model for compression. It uses a Transformer encoder to capture bidirectional contextual information, ensuring that the compressed prompt is faithful to the original. [293] proposes a technique called ’gisting’, where a language model is trained to condense the prompt into a compact ’gist token’. These tokens encapsulate the core semantic content of the prompt and can be cached for later use. This method achieves a compression rate of up to 26 times.  [102] suggests a method to transform pre-trained language models into AutoCompressors. The AutoCompressor compresses long contexts into summary vectors, and training is performed on the model parameters using these summary vectors.
@@ 2.5.3 Data Packing
Short Sequence Insertion. Some methods [116, 259] involve inserting short sequences into long sequences to minimize padding. The Best-fit Packing [116] first splits long documents according to the model’s context length, then sorts all document blocks in descending order of length. For each document block, it selects the training sequence set with the smallest remaining capacity that can accommodate it. [259] prioritizes long documents and uses a greedy algorithm to fill remaining space with short document segments (sequences), reducing padding and minimizing document concatenation to lower contextual noise.